# R3DM: Enabling Role Discovery and Diversity Through Dynamics Models in Multi-agent Reinforcement Learning

### v1

@@ 2.2 Value Function Factorization for Centralized Training with Decentralized Execution
MARL through the CTDE paradigm, such as QMIX (Rashid et al., 2020), has been a major focus in recent research due to the balance between centralized learning and decentralized decision-making. The key idea is the factorization of the value function, which composes the global action value function $Q_{\text{tot}}^{\pi}(s^{t},\mathbf{a}^{t};\phi)$$,=f(Q_{1}(\tau_{1}^{t},a_{1}^{t};\phi_{Q}),$$\dots,Q_{n}(\tau_{n}^{t},a_{n}^{t};\phi_{Q}),s^{t},\mathbf{a^{t}};\phi)$ through a function $f$, over individual utilities for each agent $Q_{i}(\tau_{i}^{t},a_{i}^{t};\phi_{Q})$ with parameters $\phi_{Q}$. By adhering to the Individual-Global-Maximum (IGM) principle, agents can independently maximize their utility functions to obtain an optimal global action, i.e., $\operatorname*{arg\,max}_{\mathbf{a}^{t}}Q_{\text{tot}}^{\pi}(s^{t},\mathbf{a}^{t};\phi)=[\operatorname*{arg\,max}_{a_{1}^{t}}Q_{1}(\tau_{1}^{t},a_{1}^{t};\phi_{Q}),...,\operatorname*{arg\,max}_{a_{n}^{t}}Q_{n}(\tau_{n}^{t},a_{n}^{t};\phi_{Q})]$.
@@ Appendix A Proofs
[…]
$\approx\mathbb{E}_{e_{i}^{t},z_{i}^{t},m_{i}^{t}}\frac{p(\tau_{i}^{t+k}\mid z_{i}^{t},\tau_{i}^{t})p(z_{i}^{t}\mid e_{i}^{t})}{p(\tau_{i}^{t+k}\mid\tau_{i}^{t})p(z_{i}^{t})}$ (We assume $(\text{We assume that}\;\;p(e_{i}^{t}\mid\tau_{i}^{t},m_{i}^{t})\approx p(e_{i}^{t}|m_{i}^{t}))$
@@
$-\mathbb{E}_{e_{i}^{t},z_{i}^{t},m_{i}^{t}}\left[\log{\frac{g(z_{i}^{t},e_{i}^{t})}{g(z_{i}^{t},e_{i}^{t})+\sum_{m_{i}^{t*}\in M/m_{i}^{t}}g(z_{i}^{t},e_{i}^{t*})}}\right]$ (13)
@@
Following the proof in  (Oord et al., 2018), it can be seen that the function in Eq. 13 is akin to the categorical loss of a model or a function $g(z_{i}^{t},e_{i}^{t})$. Therefore, $g(z_{i}^{t},e_{i}^{t})$ is the optimal probability $p(i=m_{i}^{t}|e_{i}^{t},Z)$ that the agent $i$ takes up the role $m_{i}^{t}$ based on the agent embedding $e_{i}^{t}$ and the set of all role embeddings $Z=\{z_{i}^{t}\mid i\in I\}$ for all agents in $I$
@@
$-\mathbb{E}_{e_{i}^{t},z_{i}^{t},m_{i}^{t}}\left[\log{\frac{g(z_{i}^{t},e_{i}^{t})}{g(z_{i}^{t},e_{i}^{t})+\sum_{m_{i}^{t}*\in M/m_{i}^{t}}g(z_{i}^{t},e_{i}^{t*})}}\right]\geq\log|M|-\mathbb{E}_{e_{i}^{t},z_{i}^{t},m_{i}^{t}}\left[\log\frac{p(z_{i}^{t}\mid e_{i}^{t})}{p(z_{i}^{t})}\right].$
@@ B.5 Derivation of the Intrinsic Rewards
[…]
@@ Appendix C Algorithm
12:     Append the complete trajectory to $B$ = $B\bigcup\{[o_{i}^{t}]_{i=1}^{n},s^{t},[a_{i}^{t}]_{i=1}^{n},r^{t},[e_{i}^{t}]_{i=1}^{n},[z_{i}^{t}]_{i=1}^{n}\}_{t=1}^{T}$

### v3

@@ 2.2 Value Function Factorization for Centralized Training with Decentralized Execution
MARL through the CTDE paradigm, such as QMIX (Rashid et al., 2020), has been a major focus in recent research due to the balance between centralized learning and decentralized decision-making. The key idea is the factorization of the value function, which composes the global action value function $Q_{\text{tot}}^{\pi}(s^{t},\mathbf{a}^{t};\phi)$$f(Q_{1}(\tau_{1}^{t},a_{1}^{t};\phi_{Q}),$ $\dots,Q_{n}(\tau_{n}^{t},a_{n}^{t};\phi_{Q}),s^{t},\mathbf{a^{t}};\phi)$ through a function $f$, over individual utilities for each agent $Q_{i}(\tau_{i}^{t},a_{i}^{t};\phi_{Q})$ with parameters $\phi_{Q}$. By adhering to the Individual-Global-Maximum (IGM) principle, agents can independently maximize their utility functions to obtain an optimal global action, i.e., $\operatorname*{arg\,max}_{\mathbf{a}^{t}}Q_{\text{tot}}^{\pi}(s^{t},\mathbf{a}^{t};\phi)=[\operatorname*{arg\,max}_{a_{1}^{t}}Q_{1}(\tau_{1}^{t},a_{1}^{t};\phi_{Q}),...,\operatorname*{arg\,max}_{a_{n}^{t}}Q_{n}(\tau_{n}^{t},a_{n}^{t};\phi_{Q})]$.
@@ Appendix A Proofs
$\frac{p(\tau_{i}^{t+k}|m_{i}^{t})}{p(\tau_{i}^{t+k})}$ $=\frac{\int_{e_{i}^{t},z_{i}^{t},\tau_{i}^{t}}p(\tau_{i}^{t+k}\mid z_{i}^{t},\tau_{i}^{t})p(z_{i}^{t}\mid e_{i}^{t})p(e_{i}^{t}\mid\tau_{i}^{t},m_{i}^{t})p(\tau_{i}^{t})de_{i}^{t}dz_{i}^{t}d\tau_{i}^{t}}{\int_{\tau_{i}^{t}}p(\tau_{i}^{t+k}\mid\tau_{i}^{t})p(\tau_{i}^{t})d\tau_{i}^{t}}$
$\approx\mathbb{E}_{e_{i}^{t},z_{i}^{t},m_{i}^{t}}\frac{p(\tau_{i}^{t+k}\mid z_{i}^{t},\tau_{i}^{t})p(z_{i}^{t}\mid e_{i}^{t})}{p(\tau_{i}^{t+k}\mid\tau_{i}^{t})p(z_{i}^{t})}$ $(\text{We assume that}\;\;p(e_{i}^{t}\mid\tau_{i}^{t},m_{i}^{t})\approx p(e_{i}^{t}|m_{i}^{t}))$
@@
$\mathbb{E}_{e_{i}^{t},z_{i}^{t},m_{i}^{t}}\left[\log{\frac{g(z_{i}^{t},e_{i}^{t})}{g(z_{i}^{t},e_{i}^{t})+\sum_{m_{i}^{t*}\in M/m_{i}^{t}}g(z_{i}^{t},e_{i}^{t*})}}\right]$ (13)
@@
Following the proof in  (Oord et al., 2018), it can be seen that the function in Eq. 13 is akin to the categorical loss of a model or a function $g(z_{i}^{t},e_{i}^{t})$. Therefore, $g(z_{i}^{t},e_{i}^{t})$ is the optimal probability $p(i=m_{i}^{t}|e_{i}^{t},Z)$ that the agent $i$ takes up the role $m_{i}^{t}$ based on the agent embedding $e_{i}^{t}$ and the set of all role embeddings $Z=\{z_{i}^{t}\mid i\in I\}$ for all agents in $I$ .
@@
$\mathbb{E}_{e_{i}^{t},z_{i}^{t},m_{i}^{t}}\left[\log{\frac{g(z_{i}^{t},e_{i}^{t})}{g(z_{i}^{t},e_{i}^{t})+\sum_{m_{i}^{t}*\in M/m_{i}^{t}}g(z_{i}^{t},e_{i}^{t*})}}\right]\geq\log|M|-\mathbb{E}_{e_{i}^{t},z_{i}^{t},m_{i}^{t}}\left[\log\frac{p(z_{i}^{t}\mid e_{i}^{t})}{p(z_{i}^{t})}\right].$
@@ B.5 Derivation of the Intrinsic Rewards
$\approx\mathbb{E}_{z_{i}^{t},\tau_{i}^{l},a_{i}^{l}}\left[\log\left(\frac{q_{\psi_{\mathrm{dyn}}}(d_{i}^{t+1}\mid h_{i}^{t+1},z_{i}^{t})q_{\psi_{\mathrm{dec}}}(o_{i}^{t}\mid h_{i}^{t+1},d_{i}^{t+1},z_{i}^{t})}{p_{\phi_{\mathrm{dyn}}}(d_{i}^{t+1}\mid h_{i}^{t+1})p_{\phi_{\mathrm{dec}}}(o_{i}^{t}\mid h_{i}^{t+1},d_{i}^{t+1})}\right)\right]\;\;\;\;(\text{sequence models will yield similar hidden states})$
@@ Appendix C Algorithm
12:  Append the complete trajectory to $B$ = $B\bigcup\{[o_{i}^{t}]_{i=1}^{n},s^{t},[a_{i}^{t}]_{i=1}^{n},r^{t},[e_{i}^{t}]_{i=1}^{n},[z_{i}^{t}]_{i=1}^{n}\}_{t=1}^{T}$ .